OpenStudy (anonymous):

Which of the following is an example of a recursive sequence? 2, 5, 8, 11, ... 1, 3, 9, 27, ... 2, 4, 6, 10, ...

OpenStudy (the_fizicx99):

Are you missing some info? Because a recursive sequence can also be arithmetic. Not just geometric :P

OpenStudy (anonymous):

The first is \[ a_1=2\\ a_n=a_{n-1}+3 \]

OpenStudy (anonymous):

??? @eliassaab

OpenStudy (anonymous):

5= 2+3 8=5+3; 11=8+3 What is your ????

OpenStudy (the_fizicx99):

How do you know what a_1 is O.o

OpenStudy (anonymous):

@eliassaab is it the 1st one??

OpenStudy (anonymous):

Yes it is the first one. By the way, why are you blocking me?

OpenStudy (the_fizicx99):

I would never ♥ ... 1 can also be a_1 >.>